OpenStudy (selena2345):

Evaluate The Expression 11-(-2)+14

OpenStudy (anonymous):

--2=?

OpenStudy (selena2345):

Yeah

OpenStudy (anonymous):

-(-2)=+2 or -2

OpenStudy (selena2345):

-2

OpenStudy (anonymous):

no -*-=+ 11-(-2)+14 11+ 2 +14 can you add them no

OpenStudy (selena2345):

what

OpenStudy (anonymous):

11+2=13+14=27

OpenStudy (selena2345):

Thank You My Dear